# CausalProfiler: Generating Synthetic Benchmarks for Rigorous and Transparent Evaluation of Causal Machine Learning

### v1

@@ 1 Introduction
Causal machine learning (Causal ML) seeks to estimate the effects of interventions and counterfactuals using machine learning techniques (kaddour2022causal), enabling principled decision making—for example in medicine and policy. Despite the theoretical maturity and growing relevance of Causal ML, current research practices lack rigourous evaluations of how proposed methods would perform under realistic and diverse conditions, limiting their practical utility (curth24; feuerriegel24; audrey2025; berrevoets2024causal).
In Causal ML, evaluation is particularly challenging due to the unobservability of counterfactual outcomes (holland86). Researchers can rely only on scarce real-world data sources such as randomized controlled trials, considered the gold standard, which are expensive, are ethically constrained, and often encompass a low amount of data (greenland:02; tennant:21). As a result, existing benchmarks often rely on a few semi-synthetic datasets (e.g., Syntren (syntren_06), ACIC2016 (dorie2019automated)) or model-driven synthetic datasets generated from fitted causal mechanisms (neal2020realcause; parikh2022validating; athey2024using; manela2024marginal). However, these datasets encode assumptions that are rarely made explicit and whose validity is difficult to generalize beyond the original study context (audrey2025). In parallel, many researchers define handcrafted synthetic datasets, useful for theory but fragile for empirical evaluation: a few manually chosen models can overstate performance by aligning with method-specific assumptions (gentzel2019case). Moreover, lessons from predictive machine learning show that narrow, static benchmarks can give a false sense of reliability (geirhos2020shortcut; herrmann24; freiesleben23; longjohn2024), underscoring the need for structured diversity: systematic variation of tasks under explicit, controllable assumptions.
@@
Although synthetic evaluation cannot replace real data, it offers the only reliable access to ground-truth causal queries, since counterfactuals are unobservable and many assumptions are unfalsifiable (holland86). CausalProfiler brings a much needed complement to real-world studies by enabling transparent, diverse, and controlled synthetic experiments to support method development.
@@ 2 Related Work
Evaluating Causal ML methods. Causal ML currently lacks a rigorous, systematic paradigm for empirical evaluation, whether synthetic or semi-synthetic. Semi-synthetic datasets, such as synthetic outcome datasets (dorie2019automated; shimoni2018benchmarking; hill2011bayesian) and model-based semi-synthetic datasets (neal2020realcause; parikh2022validating; athey2024using; manela2024marginal), combine real covariates and simulated outcomes under assumed structural models. On the other hand, fully synthetic datasets are generated entirely from researcher-defined SCMs, allowing for greater control and access to ground truth. Yet both synthetic and semi-synthetic approaches share critical limitations. First, synthetic evaluations often lack realism, relying on overly simplistic mechanisms such as additive noise or linear functions, and frequently omitting robustness analyses (gentzel2019case; curth24; poinsot24; audrey2025). Such evaluations rarely reflect the complexity of real-world causal processes and are insufficient to test the limits of modern causal inference methods. Secondly, synthetic and semi-synthetic datasets are shaped by researcher-defined design choices, including the causal graph structure, the form of the outcome function, and the noise distribution. These decisions, often made implicitly, can unintentionally introduce hidden biases that favor certain methods (curth2021really; cheng22; feuerriegel24). Such assumptions are rarely documented or systematically varied, hindering reproducibility and fair method comparison (poinsot24; audrey2025). Additionally, these benchmarks are typically small in scale and narrow in scope, often covering only a limited range of causal settings. As a result, empirical evaluations raise concerns about overfitting and generalization (gentzel2019case; berrevoets2024causal). For instance, it has been shown that even small changes to the data-generating process can lead to dramatic shifts in performance rankings (curth2021really). Moreover, methods are often evaluated only under the very conditions that guarantee their identifiability, offering little insight into robustness under assumption violations, as is common in real-world settings (petersen:24; hutchinson:22). In short, without broader and more transparent evaluation across diverse causal settings, the field risks drawing conclusions that do not generalize. For Causal ML to have wide impact in practice, there is a need to move beyond fixed benchmarks toward frameworks that support transparent, controlled, and diverse experimentation across well-defined spaces of causal assumptions.
@@
Recent benchmarking efforts. Recent works have sought to address some of these problems, introducing tools to generate synthetic SCMs for causal discovery (kalainathan2020causal; gupta2023learned; rudolph2023all) or support query estimation from hand-specified models (dowhy; dagitty; pymc). However, none of these frameworks support all components required for robust evaluation of causal machine learning methods. First, the causal discovery benchmarks do not compute ground truth for intervention or counterfactual queries. Further, query estimation frameworks often require manual SCM specification and do not support random sampling, diversity control, or analysis of the distribution of tasks. Even when SCMs are sampled (rudolph2023all; xia2023neural), key properties (e.g., positivity) are neither reported nor constrained. Moreover, the absence of randomness in the graph structures limits generalization. In contrast, CausalProfiler integrates SCM sampling, query ground-truth computation, and coverage guarantees into a unified framework. To the best of our knowledge, this is the first benchmark generator that enables systematic exploration of how Causal ML methods behave across spaces of SCMs and queries defined by user-specified constraints.
@@ 3 Background & Notation
We use capital letters for random variables (e.g., $X$), lowercase for realizations (e.g., $x$), and boldface for vectors (e.g., $\mathbf{x}$). For a more complete background, please refer to Appendix B and pearl09.
@@
The Pearl Causal Hierarchy (PCH) (pearl18) classifies causal reasoning into three levels: $\mathcal{L}_{1}$ (associational), $\mathcal{L}_{2}$ (interventional), and $\mathcal{L}_{3}$ (counterfactual). Associative questions use only observed data, whereas interventional and counterfactual questions require assumptions about the data-generating process. Importantly, lower levels are insufficient to answer higher-level questions in almost all causal models (bareinboim:22).
@@
The class of Structural Causal Models (SCMs) (pearl09) provide a representation allowing reasoning on the three levels of the PCH. An SCM is a tuple $\mathcal{M}:=\{\mathbf{V},\mathbf{U},\mathbfcal{F},P(\mathbf{U})\}$, where $\mathbf{V}$ is a set of endogenous variables, $\mathbf{U}$ is a set of exogenous variables, $\mathbfcal{F}$ is a set of structural equations $V_{i}=f_{i}(\textit{{PA}}(V_{i}),\mathbf{U}_{V_{i}})$, also called causal mechanisms, and $P(\mathbf{U})$ defines a distribution over the exogenous variables $\mathbf{U}$. SCMs induce a distribution $P_{\mathcal{M}}(\mathbf{V})$ over the endogenous variables $\mathbf{V}$, called the entailed distribution. We consider two types of endogenous variables: the observed variables, denoted $\mathbf{V}_{O}$, and the unobserved variables, denoted $\mathbf{V}_{H}$, where $\mathbf{V}=\mathbf{V}_{O}\cup\mathbf{V}_{H}$ and $\mathbf{V}_{O}\cap\mathbf{V}_{H}=\emptyset$. We represent causal relationships using the causal graph $\mathcal{G}$ of a SCM. This is an acyclic directed mixed graph over the endogenous variables. Directed edges $X\to Y$ encode causal dependencies via causal mechanisms where $X\in\textit{{PA}}(Y)$ is called a parent of $Y$, while bidirected edges $X\leftrightarrow Y$ indicate latent confounding due to shared exogenous causes. With SCMs one can represent intervention and counterfactual questions. An intervention replaces one or more structural equations to model external manipulations. A common example is a hard intervention, $\textit{{do}}(T=t)$, which fixes a variable’s value, disconnecting it from its causes. This defines a new SCM and alters the induced distribution. Counterfactual questions build on this idea: given an observed realization called the factual realization, they ask what would have happened under an intervention different from the one actually taken. They are evaluated by conditioning on observed variables (abduction), modifying the SCM with the intervention (action), and predicting outcomes under the new distribution (prediction)—a process known as the three-step procedure (pearl09).
@@
More generally, a causal query refers to a probabilistic statement about the effect of hypothetical manipulations of the data-generating process. This includes intervention queries, such as Average Treatment Effect (ATE), and counterfactual queries, such as Counterfactual Total Effect (Ctf-TE) (plecko:24). A query is identifiable if its value can be uniquely determined from data, given a set of assumptions (e.g., a causal sufficiency) (pearl09). In other words, identifiability refers to whether causal queries can be empirically estimated, and under what assumptions.
@@ 5.2 Sampling Structural Causal Models
Causal Graphs. CausalProfiler first samples a directed acyclic graph over a set of endogenous variables, defining the SCM’s causal structure. If specified in the SoI, CausalProfiler samples a subset of endogenous variables, $\mathbf{V}_{H}$, to be treated as unobserved and excluded from the observed dataset. To expose only the visible causal structure to the user, we apply Verma’s latent projection algorithm (verma93) to the full causal graph, which produces an acyclic directed mixed graph.
@@ 5.3 Sampling Causal Datasets
Query ground truth $Q^{\star}$. Each query is estimated by drawing samples from the (manipulated) ground truth SCM: interventional queries via do-operations (action and prediction), and counterfactual queries via the three-step procedure (pearl09).
@@
Benchmark Design. Taken together, these design choices reflect four key properties that are considered essential for rigorous synthetic evaluation in Causal ML (audrey2025): transparency, by making all assumptions explicit via the parametrization of the SoI, which serves as a declarative specification of the evaluation domain; repeatability, through randomized but seed-controlled sampling procedures, ensuring that SCMs and queries can be exactly reproduced across runs; bias awareness, supported by the coverage guarantee and the empirical distribution analysis module; and control over experiments, by exposing a wide range of configurable parameters in the SoI that allow users to tailor the causal dataset generation to their assumptions and research goals.
@@ 6.1 Verification of Benchmark Correctness
To validate the soundness of our benchmark generator, we perform consistency checks across the three levels of the PCH. Using the SCM sampler and query estimator of the CausalProfiler, we test whether sampled SCMs satisfy the Markov condition, do-calculus rules, and the structural counterfactual axioms (pearl09). We use discrete SCMs to allow exhaustive enumeration of conditioning sets for statistical tests. To ensure robustness, we iterate over a SoI parameter grid spanning the number of variables, edge density, cardinalities, and noise regions. See Appendix K for full details and results.
L1: Markov Property Verification. We test whether d-separations in the causal graph imply conditional independencies in the entailed observational distribution of the sampled SCMs. For each SCM, we enumerate d-separated triplets $(A,B,C)$ and test $A\perp B\mid C$ with Pearson’s $\chi^{2}$ test (pearson1900), filtering low-sample strata (koehler_80) and correcting for multiple tests (benjamini1995controlling). The Markov property holds in about 95% of the tested cases, with most violations due to finite-sample variability.
L2: Do-Calculus Verification. We test whether the three rules of do-calculus hold empirically. For each rule, we identify variable tuples satisfying its graphical preconditions. We then use the query estimator to generate two interventional datasets corresponding to the rule’s left- and right-hand sides. We compare the resulting distributions with Pearson’s $\chi^{2}$ test, filtering low-sample strata (koehler_80) and correcting for multiple tests (benjamini1995controlling). About 5.5% of tests fail, mostly due to finite-sample noise.
@@ 6.2 Comparison to existing benchmarks
Comparison. To illustrate CausalProfiler’s contribution to SCM diversity for evaluating Causal ML methods, we compare its SCMs (sampled over a SoI grid spanning number of variables, edge density, cardinalities, noise regions, and dataset size) with two existing benchmarks: the synthetic SCMs from the Causal Normalizing Flows (CausalNF) work (javaloy2023causal) and the CANCER and EARTHQUAKE models from bnlearn (bnlearn). For interpretable visualization, we apply two-dimensional t-SNE (maaten2008visualizing) to the computable metrics of the analysis module (Appendix G), with a perplexity set to 30, see Figure˜1.
@@ 6.3 Method Evaluation using CausalProfiler
Performance is assessed by mean squared error between predicted and true query values, with mean error, standard deviation, runtime, and failure rate (due to numerical issues or exceptions) for each method and SoI. We compare Causal Normalizing Flows (CausalNF) (javaloy2023causal), Neural Causal Models (NCM) (xia2023neural), Variational Causal Graph Autoencoder (VACA) (sanchez2021vaca), and Diffusion-based Causal Models (DCM) (chao2023interventional).
@@ 6.4 Experiment 1: General Evaluation across Diverse SCMs
Findings (Linear-Medium vs. NN-Medium). In the Linear-Medium setting, DCM achieves the lowest average error (0.1530), indicating excellent performance, but its error standard deviation is notably high (1.5289), driven by a few extreme outliers (max error 33.98). This suggests DCM is effective for most queries but can produce large errors in rare cases—potentially problematic in safety-critical applications matching this SoI. VACA performs competitively with lower max error and faster runtime, but suffers a high failure rate (53.4%) due to NaNs. In the NN-Medium setting, where the causal mechanisms are small neural networks, DCM’s advantage disappears. VACA emerges as the best performer, with the lowest error mean (0.0090) and standard deviation (0.0077), while reducing its failure rate to 5%. Interestingly, DCM becomes the weakest performer in this setting, showing that method rankings are highly sensitive to the underlying functional form of the mechanisms. This underscores the need for practitioners to evaluate methods within the SoI most relevant to their application. Lastly, NN SCMs surprisingly yield lower errors than linear ones. A plausible explanation is an inductive-bias match with the evaluated neural methods and the tendency of small randomly initialized NNs to produce relatively smooth, low-frequency functions that are easier to estimate from finite data (rahaman2019spectral).

### v2

@@ 1 Introduction
Causal machine learning (Causal ML) seeks to estimate the effects of interventions and counterfactuals using machine learning techniques (Kaddour et al., 2022), enabling principled decision making—for example in medicine and policy. Despite the theoretical maturity and growing relevance of Causal ML, current research practices lack rigourous evaluations of how proposed methods would perform under realistic and diverse conditions, limiting their practical utility (Curth et al., 2024; Feuerriegel et al., 2024; Poinsot et al., 2025; Berrevoets et al., 2024).
In Causal ML, evaluation is particularly challenging due to the unobservability of counterfactual outcomes (Holland, 1986). Researchers can rely only on scarce real-world data sources such as randomized controlled trials, considered the gold standard, which are expensive, are ethically constrained, and often encompass a low amount of data (Greenland and Brumback, 2002; Tennant et al., 2021). As a result, existing benchmarks often rely on a few semi-synthetic datasets (e.g., Syntren (den Bulcke et al., 2006), ACIC2016 (Dorie et al., 2019)) or model-driven synthetic datasets generated from fitted causal mechanisms (Neal et al., 2020; Parikh et al., 2022; Athey et al., 2024; de Vassimon Manela et al., 2024). However, these datasets encode assumptions that are rarely made explicit and whose validity is difficult to generalize beyond the original study context (Poinsot et al., 2025). In parallel, many researchers define handcrafted synthetic datasets, useful for theory but fragile for empirical evaluation: a few manually chosen models can overstate performance by aligning with method-specific assumptions (Gentzel et al., 2019). Moreover, lessons from predictive machine learning show that narrow, static benchmarks can give a false sense of reliability (Geirhos et al., 2020; Herrmann et al., 2024; Freiesleben and Grote, 2023; Longjohn et al., 2024), underscoring the need for structured diversity: systematic variation of tasks under explicit, controllable assumptions.
@@
Although synthetic evaluation cannot replace real data, it offers the only reliable access to ground-truth causal queries, since counterfactuals are unobservable and many assumptions are unfalsifiable (Holland, 1986). CausalProfiler brings a much needed complement to real-world studies by enabling transparent, diverse, and controlled synthetic experiments to support method development.
@@ 2 Related Work
Evaluating Causal ML methods. Causal ML currently lacks a rigorous, systematic paradigm for empirical evaluation, whether synthetic or semi-synthetic. Semi-synthetic datasets, such as synthetic outcome datasets (Dorie et al., 2019; Shimoni et al., 2018; Hill, 2011) and model-based semi-synthetic datasets (Neal et al., 2020; Parikh et al., 2022; Athey et al., 2024; de Vassimon Manela et al., 2024), combine real covariates and simulated outcomes under assumed structural models. On the other hand, fully synthetic datasets are generated entirely from researcher-defined SCMs, allowing for greater control and access to ground truth. Yet both synthetic and semi-synthetic approaches share critical limitations. First, synthetic evaluations often lack realism, relying on overly simplistic mechanisms such as additive noise or linear functions, and frequently omitting robustness analyses (Gentzel et al., 2019; Curth et al., 2024; Poinsot et al., 2024; 2025). Such evaluations rarely reflect the complexity of real-world causal processes and are insufficient to test the limits of modern causal inference methods. Secondly, synthetic and semi-synthetic datasets are shaped by researcher-defined design choices, including the causal graph structure, the form of the outcome function, and the noise distribution. These decisions, often made implicitly, can unintentionally introduce hidden biases that favor certain methods (Curth et al., 2021; Cheng et al., 2022; Feuerriegel et al., 2024). Such assumptions are rarely documented or systematically varied, hindering reproducibility and fair method comparison (Poinsot et al., 2024; 2025). Additionally, these benchmarks are typically small in scale and narrow in scope, often covering only a limited range of causal settings. As a result, empirical evaluations raise concerns about overfitting and generalization (Gentzel et al., 2019; Berrevoets et al., 2024). For instance, it has been shown that even small changes to the data-generating process can lead to dramatic shifts in performance rankings (Curth et al., 2021). Moreover, methods are often evaluated only under the very conditions that guarantee their identifiability, offering little insight into robustness under assumption violations, as is common in real-world settings (Petersen, 2024; Hutchinson et al., 2022). In short, without broader and more transparent evaluation across diverse causal settings, the field risks drawing conclusions that do not generalize. For Causal ML to have wide impact in practice, there is a need to move beyond fixed benchmarks toward frameworks that support transparent, controlled, and diverse experimentation across well-defined spaces of causal assumptions.
@@
Recent benchmarking efforts. Recent works have sought to address some of these problems, introducing tools to generate synthetic SCMs for causal discovery (Kalainathan et al., 2020; Gupta et al., 2023; Rudolph et al., 2023) or support query estimation from hand-specified models (Sharma and Kiciman, 2020; Textor et al., 2017; Abril-Pla et al., 2023). However, none of these frameworks support all components required for robust evaluation of causal machine learning methods. First, the causal discovery benchmarks do not compute ground truth for intervention or counterfactual queries. Further, query estimation frameworks often require manual SCM specification and do not support random sampling, diversity control, or analysis of the distribution of tasks. Even when SCMs are sampled (Rudolph et al., 2023; Xia et al., 2023), key properties (e.g., positivity) are neither reported nor constrained. Moreover, the absence of randomness in the graph structures limits generalization. In contrast, CausalProfiler integrates SCM sampling, query ground-truth computation, and coverage guarantees into a unified framework. To the best of our knowledge, this is the first benchmark generator that enables systematic exploration of how Causal ML methods behave across spaces of SCMs and queries defined by user-specified constraints.
@@ 3 Background & Notation
We use capital letters for random variables (e.g., $X$), lowercase for realizations (e.g., $x$), and boldface for vectors (e.g., $\mathbf{x}$). For a more complete background, please refer to Appendix B and Pearl (2009).
@@
The Pearl Causal Hierarchy (PCH) (Pearl and Mackenzie, 2018) classifies causal reasoning into three levels: $\mathcal{L}_{1}$ (associational), $\mathcal{L}_{2}$ (interventional), and $\mathcal{L}_{3}$ (counterfactual). Associative questions use only observed data, whereas interventional and counterfactual questions require assumptions about the data-generating process. Importantly, lower levels are insufficient to answer higher-level questions in almost all causal models (Bareinboim et al., 2022).
@@
The class of Structural Causal Models (SCMs) (Pearl, 2009) provide a representation allowing reasoning on the three levels of the PCH. An SCM is a tuple $\mathcal{M}:=\{\mathbf{V},\mathbf{U},\mathbfcal{F},P(\mathbf{U})\}$, where $\mathbf{V}$ is a set of endogenous variables, $\mathbf{U}$ is a set of exogenous variables, $\mathbfcal{F}$ is a set of structural equations $V_{i}=f_{i}(\textit{{PA}}(V_{i}),\mathbf{U}_{V_{i}})$, also called causal mechanisms, and $P(\mathbf{U})$ defines a distribution over the exogenous variables $\mathbf{U}$. SCMs induce a distribution $P_{\mathcal{M}}(\mathbf{V})$ over the endogenous variables $\mathbf{V}$, called the entailed distribution. We consider two types of endogenous variables: the observed variables, denoted $\mathbf{V}_{O}$, and the unobserved variables, denoted $\mathbf{V}_{H}$, where $\mathbf{V}=\mathbf{V}_{O}\cup\mathbf{V}_{H}$ and $\mathbf{V}_{O}\cap\mathbf{V}_{H}=\emptyset$. We represent causal relationships using the causal graph $\mathcal{G}$ of a SCM. This is an acyclic directed mixed graph over the endogenous variables. Directed edges $X\to Y$ encode causal dependencies via causal mechanisms where $X\in\textit{{PA}}(Y)$ is called a parent of $Y$, while bidirected edges $X\leftrightarrow Y$ indicate latent confounding due to shared exogenous causes. With SCMs one can represent intervention and counterfactual questions. An intervention replaces one or more structural equations to model external manipulations. A common example is a hard intervention, $\textit{{do}}(T=t)$, which fixes a variable’s value, disconnecting it from its causes. This defines a new SCM and alters the induced distribution. Counterfactual questions build on this idea: given an observed realization called the factual realization, they ask what would have happened under an intervention different from the one actually taken. They are evaluated by conditioning on observed variables (abduction), modifying the SCM with the intervention (action), and predicting outcomes under the new distribution (prediction)—a process known as the three-step procedure (Pearl, 2009).
@@
More generally, a causal query refers to a probabilistic statement about the effect of hypothetical manipulations of the data-generating process. This includes intervention queries, such as Average Treatment Effect (ATE), and counterfactual queries, such as Counterfactual Total Effect (Ctf-TE) (Plečko and Bareinboim, 2024). A query is identifiable if its value can be uniquely determined from data, given a set of assumptions (e.g., a causal sufficiency) (Pearl, 2009). In other words, identifiability refers to whether causal queries can be empirically estimated, and under what assumptions.
@@ 5.2 Sampling Structural Causal Models
Causal Graphs. CausalProfiler first samples a directed acyclic graph over a set of endogenous variables, defining the SCM’s causal structure. If specified in the SoI, CausalProfiler samples a subset of endogenous variables, $\mathbf{V}_{H}$, to be treated as unobserved and excluded from the observed dataset. To expose only the visible causal structure to the user, we apply Verma’s latent projection algorithm (Verma, 1993) to the full causal graph, which produces an acyclic directed mixed graph.
@@ 5.3 Sampling Causal Datasets
Query ground truth $Q^{\star}$. Each query is estimated by drawing samples from the (manipulated) ground truth SCM: interventional queries via do-operations (action and prediction), and counterfactual queries via the three-step procedure (Pearl, 2009).
@@
Benchmark Design. Taken together, these design choices reflect four key properties that are considered essential for rigorous synthetic evaluation in Causal ML (Poinsot et al., 2025): transparency, by making all assumptions explicit via the parametrization of the SoI, which serves as a declarative specification of the evaluation domain; repeatability, through randomized but seed-controlled sampling procedures, ensuring that SCMs and queries can be exactly reproduced across runs; bias awareness, supported by the coverage guarantee and the empirical distribution analysis module; and control over experiments, by exposing a wide range of configurable parameters in the SoI that allow users to tailor the causal dataset generation to their assumptions and research goals.
@@ 6.1 Verification of Benchmark Correctness
To validate the soundness of our benchmark generator, we perform consistency checks across the three levels of the PCH. Using the SCM sampler and query estimator of the CausalProfiler, we test whether sampled SCMs satisfy the Markov condition, do-calculus rules, and the structural counterfactual axioms (Pearl, 2009). We use discrete SCMs to allow exhaustive enumeration of conditioning sets for statistical tests. To ensure robustness, we iterate over a SoI parameter grid spanning the number of variables, edge density, cardinalities, and noise regions. See Appendix K for full details and results.
L1: Markov Property Verification. We test whether d-separations in the causal graph imply conditional independencies in the entailed observational distribution of the sampled SCMs. For each SCM, we enumerate d-separated triplets $(A,B,C)$ and test $A\perp B\mid C$ with Pearson’s $\chi^{2}$ test (Pearson, 1900), filtering low-sample strata (Koehler and Larntz, 1980) and correcting for multiple tests (Benjamini and Hochberg, 1995). The Markov property holds in about 95% of the tested cases, with most violations due to finite-sample variability.
L2: Do-Calculus Verification. We test whether the three rules of do-calculus hold empirically. For each rule, we identify variable tuples satisfying its graphical preconditions. We then use the query estimator to generate two interventional datasets corresponding to the rule’s left- and right-hand sides. We compare the resulting distributions with Pearson’s $\chi^{2}$ test, filtering low-sample strata (Koehler and Larntz, 1980) and correcting for multiple tests (Benjamini and Hochberg, 1995). About 5.5% of tests fail, mostly due to finite-sample noise.
@@ 6.2 Comparison to existing benchmarks
Comparison. To illustrate CausalProfiler’s contribution to SCM diversity for evaluating Causal ML methods, we compare its SCMs (sampled over a SoI grid spanning number of variables, edge density, cardinalities, noise regions, and dataset size) with two existing benchmarks: the synthetic SCMs from the Causal Normalizing Flows (CausalNF) work (Javaloy et al., 2023) and the CANCER and EARTHQUAKE models from bnlearn (Scutari, 2019). For interpretable visualization, we apply two-dimensional t-SNE (Maaten and Hinton, 2008) to the computable metrics of the analysis module (Appendix G), with a perplexity set to 30, see Figure˜1.
@@ 6.3 Method Evaluation using CausalProfiler
Performance is assessed by mean squared error between predicted and true query values, with mean error, standard deviation, runtime, and failure rate (due to numerical issues or exceptions) for each method and SoI. We compare Causal Normalizing Flows (CausalNF) (Javaloy et al., 2023), Neural Causal Models (NCM) (Xia et al., 2023), Variational Causal Graph Autoencoder (VACA) (Sánchez-Martin et al., 2022), and Diffusion-based Causal Models (DCM) (Chao et al., 2023).
@@ 6.4 Experiment 1: General Evaluation across Diverse SCMs
Findings (Linear-Medium vs. NN-Medium). In the Linear-Medium setting, DCM achieves the lowest average error (0.1530), indicating excellent performance, but its error standard deviation is notably high (1.5289), driven by a few extreme outliers (max error 33.98). This suggests DCM is effective for most queries but can produce large errors in rare cases—potentially problematic in safety-critical applications matching this SoI. VACA performs competitively with lower max error and faster runtime, but suffers a high failure rate (53.4%) due to NaNs. In the NN-Medium setting, where the causal mechanisms are small neural networks, DCM’s advantage disappears. VACA emerges as the best performer, with the lowest error mean (0.0090) and standard deviation (0.0077), while reducing its failure rate to 5%. Interestingly, DCM becomes the weakest performer in this setting, showing that method rankings are highly sensitive to the underlying functional form of the mechanisms. This underscores the need for practitioners to evaluate methods within the SoI most relevant to their application. Lastly, NN SCMs surprisingly yield lower errors than linear ones. A plausible explanation is an inductive-bias match with the evaluated neural methods and the tendency of small randomly initialized NNs to produce relatively smooth, low-frequency functions that are easier to estimate from finite data (Rahaman et al., 2019).